OpenStudy (anonymous):

[Algebra 2] Can anyone help explain how to do this as easy as possible? I know barely anything on this and I have a quiz on this Monday. Any help is greatly appreciated, and feel free to use any example of a question.

OpenStudy (anonymous):

More pictures on questions. Also could someone explain the formula 'period'? Not exactly sure what it does.

OpenStudy (anonymous):

Final picture. Thanks again for any help!

OpenStudy (anonymous):

Forgot to add the instructions: "Using radians, find the amplitude and period of each function. Then graph."

OpenStudy (anonymous):

@dan815

OpenStudy (amistre64):

most of this just requires memorization ....

OpenStudy (amistre64):

cos is the flip of sec .. for a visual cos <=> sec, rides along the humps of cos sin <=> csc, rides along the hums of sin tan <=> cot, is the mirror of tan and shifted.

OpenStudy (anonymous):

Could you explain how I would solve a problem such as #12 in the final picture? @amistre64

OpenStudy (amistre64):

whats our amp and period?

OpenStudy (anonymous):

Unsure what those are... can you briefly explain?

OpenStudy (amistre64):

youve already worked some problems out, its on your sheets.

OpenStudy (amistre64):

how did you determine them?

OpenStudy (anonymous):

They're notes from my teacher. He didn't really explain what they actually meant.

OpenStudy (amistre64):

amp is best viewed on cos and sin function; its represents how tall they are from the midline period is the shortest distance for which the function repeats itself

OpenStudy (amistre64):

now here is where memorization comes into play sin and cos have a normal period of 2pi tan has a normal period of pi

OpenStudy (anonymous):

Makes sense, so to find the amp of #12 wouldn't it be the number beside cos, which is 4?

OpenStudy (amistre64):

the amp of cot ... yes, its the multiplier to the function. 4 cot(u) has an amp of 4

OpenStudy (anonymous):

Okay, now to solve the rest of that problem what else would I need to do?

OpenStudy (amistre64):

if you need to determine the period ... tan(u) has a normal period of pi; its complement has a normal period of pi in other words: tan(0) = tan(pi) so lets take the \(\theta\) variable; when is it equal to pi?

OpenStudy (amistre64):

O/3 = pi, when O=??

OpenStudy (anonymous):

Wouldn't 0 be equal to the amp? Really unsure.

OpenStudy (anonymous):

@amistre64 Unsure if you're notified of my responses sorry if spamming

OpenStudy (amistre64):

4cot(u) 4 is the amp now u = x/3, simply becasue its a pain to type theta cot is periodic in pi, so when does x/3 = pi?

OpenStudy (anonymous):

Whenever theta equals pi?

OpenStudy (amistre64):

pi/3 is not equal to pi

OpenStudy (amistre64):

x/3 = pi ... its a pretty basic algebra process to solve for x

OpenStudy (anonymous):

Oh... sorry wasn't thinking so 3 cancels out itself, multiplies pi by 3 and that's it?

OpenStudy (amistre64):

so when x=3pi, therefore the period of this function is 3pi

OpenStudy (anonymous):

Okay, what would have to be done next? Would we have to move onto graphing now once the period and amp is found?

OpenStudy (amistre64):

graphing is next yes, and this takes some memorization as well this is what a normal tan(x) function looks like cot(x) = 1/tan(x)

OpenStudy (anonymous):

I know there's something called the 'unit circle', and I will have that for the test. Would that be needed?

OpenStudy (anonymous):

Could you also break down the graphing please

OpenStudy (amistre64):

imnot sure how useful the unit circle is unless your trying to find special values this is comparing tan(x) to cot(x)

OpenStudy (amistre64):

notice that at pi/4 they are normally equal to 1, our amplitude changes that now to 4

OpenStudy (anonymous):

Just to make sure, amp is for the x-coordinates and period is for y?

OpenStudy (amistre64):

no, amplitude is a measure of height ... y moves up and down x is a measure of width, x moves left and right

OpenStudy (anonymous):

So amp would have to be for Y and period would need to be for X coordinates, right?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

period means: shortest interval of X, in which the function repeats itself

OpenStudy (anonymous):

Ok and could you explain why its at 1 (y coord) due to tan? And also, the asymptote is doubled by the period correct!

OpenStudy (anonymous):

If the asymptote isn't doubled by the period, then what is? Or am I wrong and there is no asymtote

OpenStudy (amistre64):

vertical asymps are periodic as well yes tan(x) = sin(x)/cos(x) = 1 ; when sin(x) = cos(x), this is only true for x=pi/4 cos(x) = cos(x)/sin(x) = 1 .... for the same reason

OpenStudy (amistre64):

the tan (and cot) amplitude can be defined by its relevance to pi/4

OpenStudy (anonymous):

Okay... That seems to make sense. Now final question, how would I do the math for something such as #3 when trying to find the period?

OpenStudy (amistre64):

lets change the variable cos(u) let u = 0 to t, when is cos(0) = cos(t) for the first time?

OpenStudy (amistre64):

what value of t, greater than 0, is cos(0) = cos(t) ??

OpenStudy (anonymous):

Sorry one quick question what determines tan/sine/cos when finding the period?

OpenStudy (amistre64):

memorization ..... trig is 98% memorization which is why its such a pain

OpenStudy (anonymous):

So what would you suggest to finding it the easiest way?

OpenStudy (amistre64):

memorize when cos(0) comes around again. cos(0) equals what?

OpenStudy (anonymous):

Unsure how to determine it with the 3 relationships sine tan and cos

OpenStudy (amistre64):

what is the function define in #3?

OpenStudy (anonymous):

Cos = sin?

OpenStudy (amistre64):

\[y=2+4cos(u)\] what function is being used in the definition of y?

OpenStudy (anonymous):

Cos?

OpenStudy (amistre64):

yeah, cos is used in the definition so we apply it to the graph

OpenStudy (amistre64):

we could just as well solve: x1 /3 + pi/2 = 0 x2 /3 + pi/2 = 2pi and then draw the cosine along the interval form x1 to x2; for a period of x2-x1

OpenStudy (anonymous):

Oh.. The cosine part's answer was right there... One thing though when simplifying did you have 2 times itself and times the entire equation? How would you have gotten x2-x1 more specifically?

OpenStudy (anonymous):

Also what did you do with the period to turn it into x6 according to the graph?

OpenStudy (amistre64):

the argument for the function, the stuff thats inside the (....) it modifies the x axis. factor out the coefficient of x u = x/3 + pi/2 = 1/3 (x + 3pi/2) does this make sense?

OpenStudy (amistre64):

2 +- Amp is the range of the graphs height for #3 amp = 4 soo 2 +- 4 = -2 to 6

OpenStudy (anonymous):

Somewhat, but doesn't fully explain how you got x6 on the graph as your period

OpenStudy (amistre64):

edit **** the amplitude ISNT x6, its 4 ^^^^^^^^^^^ the normal midline of cos is y=0 when we add 2 to it its shifts the graph upwards by ... well 2 y = 2 + 4cos(u); cos(u) = -1 to 1 soo max y = 2+4 = 6 min y = 2-4 = -2 there is no x6 amplitude.

OpenStudy (anonymous):

I get what you mean as for amp now, it can go as low as -2 to as high as 6 correct?

OpenStudy (anonymous):

Sorry I meant 6pi for period

OpenStudy (amistre64):

the range of y is therefore -2 to 6, a distance of 8; and 8/2 = 4 ... our amp so your question is how to determine the period which is why i asked the question earlier: cos(0) = cos(t) for what value of t?

OpenStudy (anonymous):

When you were solving for period it didn't really say how you got 6pi, its the only thing I don't get at this point

OpenStudy (amistre64):

x1 /3 + pi/2 = 3(0-pi/2); when x1 =-3pi/2 x2 /3 + pi/2 = 3(2pi - pi/2); when x2 = 9pi/2 the interval of X that gives us ONE copy of this function is therefore from: -3pi/2 to 9pi/2 the distance between x1 and x2 is 9pi/2 - -3pi/2 = 12pi/2 the period is therefore 12pi/2, or 6pi

OpenStudy (anonymous):

Thanks for the help.... Your latest response really cleared that up. Will you be on tomorrow in case I were to run into any problems with graphing?

OpenStudy (amistre64):

the shortcut to period ... is to take the coefficient of x, and divide 2pi by it x/3 has a coeff of 1/3 2pi --- = 3*2pi = 6pi 1/3

OpenStudy (amistre64):

i cant guarentee what im doing in 3 seconds much less 6 hours :)

OpenStudy (amistre64):

and i seem to have let my brain wander .. that calculations are: x1 /3 + pi/2 = 0 x1 /3 = 0-pi/2 x1 =-3pi/2 x2 /3 + pi/2 = 2pi x2 /3 = 2pi - pi/2 x2 = 3(2pi - pi/2) x2 = 9pi/2

OpenStudy (anonymous):

Great help, I have a decent chance of doing well now. Thanks!

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

@amistre64 looking at the way you did the period, could you break this down in the easiest way you know how and explain where the values came from? I want to make sure I understand it completely and one more thing could you explain how you would find the period if you had fractions in the equation adding/subtracting with theta involved?

OpenStudy (amistre64):

when does a circle, complete itself and start over again? trig is all about circle function ... and relating them to right triangles. the unit circle is a measure of how many radians?

OpenStudy (amistre64):

and lets call theta t instead

OpenStudy (anonymous):

A circle completes itself when going all the way around itself, which is 360 degrees; the unit circle would be 360 in radians?

OpenStudy (amistre64):

yes, now lets place the unit circle so that 0 degree is the positivie x axis, and work it up and over; let x=cos(t) and y=sin(t)  the period of sin(t) and cos(t) is one complete circle; they start their values over again after one complete turn. memorize; period of sin and cos = 2pi, or 360 degrees